SuicidalSheep32:

As shown below, an observer (O) is located 660 feet from a tree (T). The observer notices a hawk (H) flying at a 35° angle of elevation from his line of sight. What equation and trigonometric function can be used to solve for the height (h) of the hawk? What is the height of the hawk? You must show all work and calculations to receive full credit.

SuicidalSheep32:

1 attachment
Gdeinward:

A: Tanjent35=x/660 ft. We do this because to find the height we must use tangent and the adjacent side to angle 35. Tanjent35=x/660 Tangent35 x 660 = x (multiply each side by 360 to isolate our variable) Tangent of 35 =0.7002075382( by using a scientific calculater) 0.7002075382 x 660 = 462.1369752 462.1369752=x The height of the bird is 462.14 ft.